# Damien Hutsemékers    Jean Manfroid    Emmanuël Jehin    Cyrielle Opitom    Michele Bannister    Juan Pablo Carvajal    Rosemary Dorsey    K Aravind    Baltasar Luco    Brian Murphy    Thomas H. Puzia    Rohan Rahatgaonkar

### v1

@@
11institutetext: Institut d’Astrophysique et de Géophysique, Université de Liège, Allée du 6 Août 19c, 4000 Liège, Belgium 22institutetext: Institute for Astronomy, University of Edinburgh, Royal Observatory, Edinburgh EH9 3HJ, United Kingdom 33institutetext: School of Physical and Chemical Sciences - Te Kura Matū, University of Canterbury, Private Bag 4800, Christchurch 8140, New Zealand 44institutetext: Institute of Astrophysics, Pontificia Universidad Católica de Chile, Av. Vicuña Mackenna 4860, 7820436 Macul, Santiago, Chile 55institutetext: Department of Physics, University of Helsinki, P.O. Box 64, 00014 Helsinki, Finland
Extreme NiI/FeI abundance ratio in the coma of the interstellar comet 3I/ATLASthanks: Based on observations made with the ESO Very Large Telescope at the Paranal Observatory under programs 115.C-0282 and 115.C-2317.
Damien Hutsemékers    Jean Manfroid    Emmanuël Jehin    Cyrielle Opitom    Michele Bannister    Juan Pablo Carvajal    Rosemary Dorsey    K Aravind    Baltasar Luco    Brian Murphy    Thomas H. Puzia    Rohan Rahatgaonkar
@@
Emission lines of FeI and NiI are commonly found in the coma of solar system comets, even at large heliocentric distances. These atoms are most likely released from the surface of the comet’s nucleus or from a short-lived parent. The presence of these lines in cometary spectra is unexpected because the surface blackbody equilibrium temperature is too low to allow the sublimation of refractory minerals containing these metals. These lines were also found in the interstellar comet 2I/Borisov which has a NiI/FeI abundance ratio similar to that observed in solar system comets. On average, this ratio is one order of magnitude higher than the solar Ni/Fe abundance ratio. Here, we report observations of the new interstellar comet 3I/ATLAS, which were carried out with the ESO Very Large Telescope equipped with the UVES spectrograph. Spectra were obtained at six epochs, at heliocentric distances ranging from 3.14 to 2.14 au. NiI was detected at all epochs. FeI was only detected at heliocentric distances smaller than 2.64 au. We estimated the NiI and FeI production rates by comparing the observed line intensities with those produced by a fluorescence model. Comet 3I exhibits a high production rate of NiI atoms as well as a high NiI/FeI ratio, making it exceptional when compared to solar system comets and 2I/Borisov. Additionally, we found that the NiI/FeI ratio decreases rapidly with decreasing heliocentric distance, suggesting that comet 3I could soon become indistinguishable from solar system comets in this respect. We interpreted these observations assuming that the NiI and FeI atoms were released through the sublimation of Ni(CO)4 and Fe(CO)5 carbonyls, which supports the presence of these species in the cometary material.
@@ 1 Introduction
Interstellar comets, with potentially different chemical compositions, offer a unique opportunity to better understand the origin of the NiI and FeI atoms observed in cometary comae. FeI and NiI were found in the interstellar comet 2I/Borisov (Guzik & Drahus 2021; Opitom et al. 2021) with a NiI/FeI abundance ratio similar to the solar system comets. NiI was recently found in the interstellar comet 3I/ATLAS at a heliocentric distance of 3.88 au, before the onset of CN and FeI (Rahatgaonkar et al. 2025). This could support an unusual composition of comet 3I, as suggested by the very high CO2/H2O abundance ratio found by Cordiner et al. (2025).
In this work, we report the first detection of FeI emission lines in the coma of comet 3I/ATLAS at a heliocentric distance of 2.64 au. We measured the NiI/FeI abundance ratio at different heliocentric distances and compared it to the ratio observed in solar system comets and comet 2I/Borisov to determine whether comet 3I is unique.
2 Observations, data reduction, and measurements
Observations were carried out with the Very Large Telescope (VLT) at the European Southern Observatory (ESO), equipped with the UV-Visual Echelle Spectrograph(UVES [^1] UVES User Manual, VLT-MAN-ESO-13200-1825,
https://www.eso.org/sci/facilities/paranal/instruments/instruments.html
). The standard settings 346+580 (dichroic 1), 390+580 (dichroic 1), and 437+860 (dichroic 2) were used. The blue settings centered at 346 nm, 390 nm and 437 nm cover the spectral ranges 3030-3880 Å, 3260-4540 Å, and 3730-4990 Å, respectively. A non-standard setting centered at 348 nm, corresponding to the spectral range 3100-3900 Å, was also used to simultaneously cover the regions containing the OH and CN bands. Most of the time, a 1.8″-wide slit was used, delivering a resolving power of about 35000. The observing circumstances are summarized in Table 1.
Raw frames were first processed to remove cosmic ray hits using the python implementation of the “lacosmic” package (van Dokkum 2001; van Dokkum et al. 2012). The data were reduced with the UVES pipeline [^2] UVES Pipeline User Manual, VLT-MAN-ESO-19500-2965,
@@ 2 Observations, data reduction, and measurements
to obtain wavelength and flux calibrated two-dimensional spectra. One-dimensional spectra were then extracted by integrating over the slit length (Table 2), using custom procedures. The scattered spectrum of the Sun (dust, twilight) was removed as described in Manfroid et al. (2009).
The intensities of unblended FeI and NiI emission lines were measured in the different spectra and compared to the intensities computed with a dedicated fluorescence model to derive FeI and NiI column densities. The procedure is described in detail in Manfroid et al. (2021). To compute the production rates from the column densities, we assume that the emission line surface brightness decreases as $p^{-1}$, where $p$ is the projected nucleocentric distance, as observed in solar system comets (Manfroid et al. 2021; Hutsemékers et al. 2021), and in 3I itself (Fig. 1 and Sect. 3). The resulting production rates and abundance ratios are given in Table 2, together with the number of lines used in the analysis.
The production rates of OH, CN, and C2 were also estimated, using the OH(0-0) band at 3090 Å, the CN(0-0) band at 3870 Å, and the C2 band at 5140 Å. The rates were derived using the fluorescence efficiencies and Haser scalelengths from Schleicher & A’Hearn (1988), Cochran & Schleicher (1993), A’Hearn et al. (1995), and Schleicher (2010), with a parent and daughter velocity of 0.85 $r_{h}^{-0.5}$ km s-1 (Cochran & Schleicher 1993). These production rates are given in Table 3.
Figure 1: Spatial profile of the brightest NiI line ($\lambda$ 3458Å) observed with UVES on September 4. The measured surface brightness (SB; normalized to one at the photocenter) is plotted as a function of the projected nucleocentric distance $p$ in arcsec. The red line represents SB $\propto p^{-1}$ convolved with a 1.5″ FWHM Gaussian to account for the seeing and tracking imperfections.
3 Analysis and results
@@
Figure 2: Continuum-subtracted spectra of comet 3I obtained on August 28 and September 3+4. Three narrow FeI lines are detected on August 28 in this spectral range, FeI $\lambda$3719.93Å being the brightest iron line detected in our spectra. On the average spectrum obtained in September 3+4 (shifted vertically for clarity), the FeI lines are brighter, and four of them are detected. Broad emission features due to background [OII] emission at 3727Å and 3729Å are also observed.
@@
Figure 3: The ratio $\log_{10}(I_{\rm obs}/I_{\rm mod})$ for the FeI (red diamonds) and NiI (blue squares) lines observed in comet 3I on August 28 (top) and September 12 (bottom). $I_{\rm obs}$ represents the observed line intensities and $I_{\rm mod}$ the intensities computed with the fluorescence model. $I_{\rm obs}/I_{\rm mod}$ is proportional to the column density of the atoms (see Manfroid et al. 2021, for details). The separation between the mean values computed for NiI and FeI gives the NiI/FeI abundance ratio. The ratios have been shifted on the y-axis so that the mean value of $\log_{10}(I_{\rm obs}/I_{\rm mod})$ is zero for FeI.
@@
One important characteristic of the FeI and NiI emission lines is their short spatial extent. In spectra obtained for comets 103P/Hartley2 and 46P at geocentric distances of 0.17 and 0.09 au, respectively, the surface brightness was found to be inversely proportional to the projected distance to the nucleus (Manfroid et al. 2021; Hutsemékers et al. 2021). Such a spatial profile indicates that the FeI and NiI atoms originate at nucleocentric distances smaller than about 50 km, either ejected from the surface of the nucleus, or released from a short-lived parent. As shown in Fig. 1, a similar profile is observed for comet 3I.
Figure 4: Production rate ratio Q(FeI+NiI)/Q(H2O+CO) versus Q(CO)/Q(H2O) for interstellar and solar system comets. The CO and H2O production rates of comet 3I are from Cordiner et al. (2025), when the comet was at 3.32 au. The CO and H2O production rates of the other comets are from the compilation of Manfroid et al. (2021). The value of Q(FeI+NiI) = Q(NiI) which was considered for comet 3I is taken when the comet was at 3.14 au (August 12, see Table 2).
@@
Figure 5: Q(NiI)/Q(FeI) as a function of the heliocentric distance $r_{h}$ for interstellar and solar system comets. The measurements obtained at different heliocentric distances are not averaged and shown individually for comets 3I (4 values, pre-perihelion), 2I/Borisov (2 values, post-perihelion), C/2001 P1 (4 values, pre-perihelion), and C/2017 K2 (2 values, pre-perihelion). The comet with the smallest Q(NiI)/Q(FeI) ratio is C/1965 S1(Ikeya-Seki) observed at 0.14 au from the Sun. The solar Ni/Fe abundance ratio is equal, in logarithm, to $1.25\pm 0.04$ (Asplund et al. 2009).
@@
Figure 7: Q(NiI)/Q(FeI) as a function of the total production rate Q(FeI+NiI).
While NiI lines were observed in the spectra obtained before August 28 and detected as far as 3.88 au (Rahatgaonkar et al. 2025, see also Table 2), FeI lines were only detected in the spectra obtained on August 28 and after that date, when the comet was at $r_{h}\leq 2.64$ au. Some of the brightest FeI lines are shown in Fig. 2. After August 28, the lines steadily intensified, and new ones emerged until our final observations in mid-September (Figs. 10 and 11). The comparison of observed and computed line intensities is shown in Fig. 3, for two epochs. Although there is clearly some dispersion between the individual measurements, the average of the intensity ratios is well defined for both atomic species at the different epochs. The difference of the average values (in logarithm) gives the NiI/FeI abundance ratio (in logarithm), which is clearly higher on August 28 than on September 12 (Table 2).
We compared the NiI and FeI production rates and the NiI/FeI abundance ratio measured in comet 3I to the same quantities evaluated in other comets (see Figs. 4-7). The NiI and FeI production rates were measured in approximately 20 comets based on UVES observations, except for comet C/1996 B2 (Hyakutake). The data are taken from Manfroid et al. (2021) and Hutsemékers et al. (2021), as are the production rates of C2 and CN. These data include those obtained for the interstellar comet 2I/Borisov (Opitom et al. 2021). We also added the data of C/2017 K2 which was recently investigated by Hmiddouch et al. (2025). Unless otherwise stated, when more than one measurement is available for a given comet, we used the simple average, with errors computed from the individual measurements. For comet 3I, however, we show the individual measurements separately. Solar-system comets were classified into two broad categories: Jupiter-family comets (JFCs) and Oort-cloud comets (OCCs). For subgroup definitions and the identification of the individual comets in the plots, we refer to Manfroid et al. (2021) and Hutsemékers et al. (2021). In the plots, we emphasized the measurements of the interstellar comet 2I/Borisov and of the unusual N2-rich solar system comet C/2016 R2 (Opitom et al. 2019).
Figure 4 shows the ratio of Q(FeI+NiI) to Q(H2O+CO) as a function of the ratio of Q(CO) to Q(H2O), for the comets in the sample studied so far. Comets 3I, 2I/Borisov, and C/2016 R2, with their high CO/H2O abundance ratios, are clearly distinct from the other comets. The observed correlation indicates that the release of NiI and FeI atoms begins before H2O sublimates, when the comets are still very cold and their activity likely driven by the more volatile CO and CO2. Figs. 5 through 7 show the production rate ratio Q(NiI)/Q(FeI) as a function of various quantities. In all these plots, comet 3I shows extreme properties, while comets 2I/Borisov and C/2016 R2 follow the global trends. Fig. 5 shows Q(NiI)/Q(FeI) as a function of the heliocentric distance. No correlation was found between Q(NiI)/Q(FeI) and $r_{h}$ for the solar system comets (Manfroid et al. 2021). For some of them, Q(NiI)/Q(FeI) was measured at different heliocentric distances, and was found to be constant within the error bars. However, comet 3I breaks the rules by showing an extremely high Q(NiI)/Q(FeI) ratio at $r_{h}\simeq$ 2.64 au, and a rapid decrease with decreasing $r_{h}$. After two weeks, its value becomes compatible with that of the other solar system comets. In fact, we cannot exclude the possibility that some other comets exhibited the same behavior but they were simply not observed early enough with a sufficient signal to noise ratio.
A similar behavior is seen in Fig. 6, which shows the correlation found between Q(NiI)/Q(FeI) and Q(C2)/Q(CN) (Hutsemékers et al. 2021): the Q(NiI)/Q(FeI) ratio of comet 3I measured in the first epochs is clearly higher than that of the other comets, but it rapidly decreases to align with them. Comet 3I appears as a C2-depleted comets at all epochs, according to the criteria defined by Bair & Schleicher (2025) based on a large sample of solar-system comets. Although the errors are large for the first two epochs, Q(C2)/Q(CN) seems to decrease with decreasing heliocentric distance, which is quite unusual; in most solar system comets, Q(C2)/Q(CN) can slightly decrease but with increasing heliocentric distance, which is an artifact produced by the oversimplified Haser model used when deriving the C2 production rates (Bair & Schleicher 2025).
In Fig. 7, comet 3I shows its most extreme behavior. While some solar system comets show either a NiI+FeI production rate or a NiI/FeI abundance ratio as large as those measured in comet 3I, comet 3I is characterized by both a high production rate of NiI atoms and a high Q(NiI)/Q(FeI) ratio.
4 Discussion and conclusions
Figure 8: Q(NiI)/Q(FeI) as a function of the heliocentric distance for interstellar and solar system comets. The comets with Q(C2)/Q(CN) << 1 are represented in red, while comets with Q(C2)/Q(CN) >> 1 are represented in blue. C/1965 S1(Ikeya-Seki) (in black) has no Q(C2)/Q(CN) measurement. The curves represent the ratio of the sublimation rates of the Ni(CO)4 and Fe(CO)5 carbonyls (see text and Manfroid et al. 2021), assuming either $r_{h}=(280/T)^{2}$ (solid line) or $r_{h}=(350/T)^{2}$ (dotted line).
Figure 9: Q(NiI) in comet 3I as a function of the heliocentric distance. The data are from this work and Rahatgaonkar et al. (2025). The curve represents the sublimation rate of the Ni(CO)4 carbonyl (Manfroid et al. 2021), assuming $r_{h}=(280/T)^{2}$. The sublimation rate is computed in molecules m-2 s-1, and shifted vertically by 0.1dex to fit the observed production rates (in molecules s-1). Thus, this shift corresponds to a very small effective emitting area of about 1.3 m2, as could be expected for a minor species.

### v2

# Damien Hutsemékers    Jean Manfroid    Emmanuël Jehin    Cyrielle Opitom    Rohan Rahatgaonkar    Michele Bannister    Juan Pablo Carvajal    Rosemary Dorsey    K. Aravind    Baltasar Luco    Brian Murphy    Thomas H. Puzia

@@
11institutetext: STAR, Institut d’Astrophysique et de Géophysique, Université de Liège, Allée du 6 Août 19c, 4000 Liège, Belgium 22institutetext: Institute for Astronomy, University of Edinburgh, Royal Observatory, Edinburgh EH9 3HJ, United Kingdom 33institutetext: School of Physical and Chemical Sciences - Te Kura Matū, University of Canterbury, Private Bag 4800, Christchurch 8140, New Zealand 44institutetext: Institute of Astrophysics, Pontificia Universidad Católica de Chile, Av. Vicuña Mackenna 4860, 7820436 Macul, Santiago, Chile 55institutetext: Department of Physics, University of Helsinki, P.O. Box 64, 00014 Helsinki, Finland
Pre-perihelion evolution of the NiI/FeI abundance ratio in the coma of the interstellar comet 3I/ATLAS. From extreme to normalthanks: Based on observations made with the ESO Very Large Telescope at the Paranal Observatory under programs 115.C-0282, 115.C-2317, and 2115.C-5030.
Damien Hutsemékers    Jean Manfroid    Emmanuël Jehin    Cyrielle Opitom    Rohan Rahatgaonkar    Michele Bannister    Juan Pablo Carvajal    Rosemary Dorsey    K. Aravind    Baltasar Luco    Brian Murphy    Thomas H. Puzia
@@
Emission lines of FeI and NiI are commonly found in the coma of Solar System comets, even at large heliocentric distances. These atoms are most likely released from the surface of the comet’s nucleus or from a short-lived parent. The presence of these lines in cometary spectra is unexpected because the surface blackbody equilibrium temperature is too low to allow the sublimation of refractory minerals containing these metals. These lines were also found in the interstellar comet 2I/Borisov, which has a NiI/FeI abundance ratio similar to that observed in Solar System comets. On average, this ratio is one order of magnitude higher than the solar Ni/Fe abundance ratio. Here, we report observations of the interstellar comet 3I/ATLAS, which were carried out with the ESO Very Large Telescope equipped with the UVES and X-shooter spectrographs. Spectra were obtained at heliocentric distances ranging from 3.14 to 1.85 au. NiI was detected at all epochs. FeI was only detected at heliocentric distances smaller than 2.64 au. We estimated the NiI and FeI production rates by comparing the observed line intensities with those produced by a dedicated fluorescence model. Comet 3I first exhibited extreme and unusual NiI/FeI abundance ratios during the initial stages of its activity. However, as its heliocentric distance decreased, this ratio became indistinguishable from those observed in Solar System comets and in comet 2I/Borisov. Comet 3I was found to be C2-depleted, with a NiI/FeI abundance ratio finally consistent with other C2-depleted comets. Nevertheless, comet 3I remains exceptional due to its high, total production rate of NiI and FeI, which is at least one order of magnitude larger than that of other comets. We interpreted these observations assuming that the NiI and FeI atoms were released through the sublimation of Ni(CO)4 and Fe(CO)5 carbonyls. This scenario provides a straightforward explanation for the asymmetric release of NiI and FeI atoms in the cometary coma and how it depends on the heliocentric distance. It also supports the presence of carbonyls in the cometary material.
@@ 1 Introduction
Interstellar comets, with potentially different chemical compositions, offer a unique opportunity to better understand the origin of the NiI and FeI atoms observed in cometary comae. FeI and NiI were found in the interstellar comet 2I/Borisov (hereafter 2I) with a NiI/FeI abundance ratio similar to the Solar System comets (Guzik & Drahus 2021; Opitom et al. 2021). NiI was recently found in the interstellar comet 3I/ATLAS (hereafter 3I) at a heliocentric distance of 3.88 au, before the onset of CN and FeI (Rahatgaonkar et al. 2025). This could support an unusual composition of comet 3I, as is suggested by the very high CO2/H2O abundance ratio found by Cordiner et al. (2025).
In this work, we report the first detection of FeI emission lines in the coma of comet 3I at a heliocentric distance of 2.64 au. We measured the NiI/FeI abundance ratio at different heliocentric distances and compared it to the ratio observed in Solar System comets and comet 2I to determine whether comet 3I is unique. We show that the release of NiI and FeI atoms at large distances from the Sun can be explained by the sublimation of short-lived species, possibly carbonyls.
2 Observations and data reduction
Observations were carried out from August 12 to September 14, 2025, with the Very Large Telescope (VLT) at the European Southern Observatory (ESO), equipped with the UV-Visual Echelle Spectrograph (UVES [^1] UVES User Manual, VLT-MAN-ESO-13200-1825,
https://www.eso.org/sci/facilities/paranal/instruments.html
). The standard settings 346+580 (dichroic 1), 390+580 (dichroic 1), and 437+860 (dichroic 2) were used. The blue settings centered at 346 nm, 390 nm, and 437 nm cover the spectral ranges 3030-3880 Å, 3260-4540 Å, and 3730-4990 Å, respectively. A nonstandard setting centered at 348 nm, corresponding to the spectral range 3100-3900 Å, was also used to simultaneously cover the regions containing the OH and CN bands. Most of the time, a 1.8″-wide slit was used, delivering a resolving power of about 35000. The observing circumstances are summarized in Table 1. Raw frames were first processed to remove cosmic ray hits using the Python implementation of the “lacosmic” package (van Dokkum 2001; van Dokkum et al. 2012). The data were reduced with the UVES pipeline [^2] UVES Pipeline User Manual, VLT-MAN-ESO-19500-2965,
@@ 2 Observations and data reduction
to obtain wavelength- and flux-calibrated two-dimensional spectra. One-dimensional spectra were then extracted by integrating over the full slit length (Table 3), using custom procedures. The scattered spectrum of the Sun (dust, twilight) was removed in the manner described in Manfroid et al. (2009).
We also considered the spectra obtained with the VLT X-shooter spectrograph [^3] X-shooter User Manual, ESO-270545,
https://www.eso.org/sci/facilities/paranal/instruments.html
(Vernet et al. 2011) and described in detail by Rahatgaonkar et al. (2025), in particular the UVB spectra obtained from July 23 to August 21, 2025, which contain the first detections of NiI lines in comet 3I. These observations were secured with a 1.6″-wide slit, providing a resolving power of about 3200. Another series of spectra, containing both the NiI and FeI lines, was obtained from September 20 to 25, 2025, with the same instrumental settings. The observing circumstances are summarized in Table 2. All of these spectra were independently reduced in the present work. The X-shooter pipeline [^4] X-shooter Pipeline User Manual, VLT-MAN-ESO-14650-4840,
https://www.eso.org/sci/software.html
was used to obtain wavelength- and flux- calibrated two-dimensional spectra. Since the target was not always positioned in the middle of the 10″-long slit, the one-dimensional spectra were extracted over a shorter, 7.5″-long slit centered on the target. The scattered spectrum of the Sun was subtracted using the solar analog star HD150469, which was observed with the same settings as the comet (Rahatgaonkar et al. 2025).
3 FeI line detection and production rate measurements
@@
Figure 1: Continuum-subtracted spectra of comet 3I obtained on August 28 and September 3+4 with UVES. Three narrow FeI lines were detected on August 28 in this spectral range, FeI $\lambda$3719.93Å being the brightest iron line detected in our spectra. In the average spectrum obtained in September 3+4 (shifted vertically for clarity), the FeI lines are brighter, and four of them were detected (as is indicated by the vertical red lines). Broad emission features due to background [OII] emission at 3727Å and 3729Å are also observed.
@@
Figure 2: Ratio $\log_{10}(I_{\rm obs}/I_{\rm mod})$ for the FeI (red diamonds) and NiI (blue squares) lines measured in comet 3I for observations secured with UVES on August 28 (top) and September 12 (bottom). $I_{\rm obs}$ represents the observed line intensities and $I_{\rm mod}$ the intensities computed with the fluorescence model. $I_{\rm obs}/I_{\rm mod}$ is proportional to the column density of the atoms (Manfroid et al. 2021). The separation between the mean values computed for NiI and FeI gives the NiI/FeI abundance ratio. The ratios have been shifted on the y axis so that the mean value of $\log_{10}(I_{\rm obs}/I_{\rm mod})$ is zero for FeI.
Figure 3: Spatial profile of the brightest NiI line ($\lambda$ 3458Å) observed with UVES on September 4. The measured surface brightness (SB; normalized to one at the photocenter) is plotted as a function of the projected nucleocentric distance, $p$, in arcsec. The red line represents SB $\propto p^{-1}$ convolved with a 1.5″ full-width-at-half-maximum Gaussian to account for the seeing and tracking imperfections.
While NiI lines were observed in the spectra obtained before August 28 and detected as far as 3.88 au (Rahatgaonkar et al. 2025, see also Table 3), FeI lines were only detected in the spectra obtained on August 28 with UVES, and after that date when the comet was at $r_{h}\leq 2.64$ au. Some of the brightest FeI lines are shown in Fig. 1. After August 28, the lines steadily intensified, and new ones emerged (Figs. 9 and 10).
The intensities of unblended FeI and NiI emission lines were measured in the UVES and X-shooter spectra for each epoch, and compared to the intensities computed with a dedicated fluorescence model to derive the FeI and NiI column densities. This model, described in detail in Manfroid et al. (2021), considers a large number of transitions and atomic levels for both the FeI and NiI atoms. Since FeI and NiI lines are also present in absorption in the solar light that irradiates the cometary atmosphere, the excitation can strongly depend on the Doppler shift between the comet and the solar spectra. Therefore, the high-resolution structure of the solar spectrum and the relative velocity between the comet and the Sun must be considered.
A comparison of observed and computed line intensities for FeI and NiI is shown in Fig. 2, for two epochs. The ratio $I_{\rm obs}/I_{\rm mod}$ gives the atomic column density (Manfroid et al. 2021). Although there is clearly some dispersion between the individual measurements, the average of the intensity ratios is well defined for both atomic species. The difference of the average values gives the NiI/FeI abundance ratio, which is clearly higher on August 28 than on September 12 (see also Table 3).
@@
One important characteristic of the FeI and NiI emission lines is their short spatial extent. In spectra obtained for comets 103P/Hartley2 and 46P at geocentric distances of 0.17 and 0.09 au, respectively, the surface brightness was found to be inversely proportional to the projected distance from the nucleus, $p$ (Manfroid et al. 2021; Hutsemékers et al. 2021). Such a spatial profile indicates that the FeI and NiI atoms originate at small nucleocentric distances, either directly ejected from the surface of the nucleus or released from a short-lived parent. As is shown in Fig. 3, a similar profile is observed for comet 3I.
Assuming that the emission line surface brightness decreases as $p^{-1}$ for both NiI and FeI, production rates were computed from the column densities according to the formulae given in Manfroid et al. (2021). We assumed a constant ejection velocity, $v$ = 0.85 $r_{h}^{-0.5}$ km s-1 (Cochran & Schleicher 1993). The resulting production rates and abundance ratios are given in Table 3, together with the number of lines used in the analysis. The NiI production rates measured for the first X-shooter dataset are in good agreement with those reported in Rahatgaonkar et al. (2025), although they are slightly smaller due a more conservative measurement of the line intensities. Since NiI and FeI were simultaneously observed, abundance ratios can be derived from either column densities or production rates.
The production rates of OH, CN, and C2 were also estimated as in Manfroid et al. (2021), using the OH(0-0) band at 3090 Å, the CN(0-0) band at 3870 Å, and the C2 band at 5140 Å. The rates were derived using the fluorescence efficiencies and Haser scalelengths from Schleicher & A’Hearn (1988), Cochran & Schleicher (1993), A’Hearn et al. (1995), and Schleicher (2010), with a parent and daughter velocity of 0.85 $r_{h}^{-0.5}$ km s-1 (Cochran & Schleicher 1993). These production rates are given in Table 4.
4 Comparison of comet 3I to other comets
@@
Figure 4: Q(NiI)/Q(FeI) as a function of the heliocentric distance, $r_{h}$, for interstellar and Solar System comets. The measurements obtained at different heliocentric distances are not averaged and shown individually for comets 3I (9 values, pre-perihelion), 2I (2 values, post-perihelion), C/2001 P1 (4 values, pre-perihelion), and C/2017 K2 (2 values, pre-perihelion). The comet with the smallest Q(NiI)/Q(FeI) ratio is C/1965 S1(Ikeya-Seki) observed at 0.14 au from the Sun. The solar Ni/Fe abundance ratio is equal, in logarithm, to $1.25\pm 0.04$ (Asplund et al. 2009).
We compared the NiI and FeI production rates and the NiI/FeI abundance ratio measured in comet 3I to the same quantities evaluated in other comets (Figs. 4 to 6). The NiI and FeI production rates were measured in approximately 20 comets based on UVES observations (except comet C/1996 B2 Hyakutake). The data are taken from Manfroid et al. (2021) and Hutsemékers et al. (2021), as are the production rates of C2 and CN. These measurements include those obtained for the interstellar comet 2I (Opitom et al. 2021). We also added the data of comet C/2017 K2, which was recently investigated by Hmiddouch et al. (2025). Unless otherwise stated, when more than one measurement is available for a given comet, we used the simple average, with errors computed from the individual measurements. For comet 3I, however, we show the individual measurements separately. Solar System comets were classified into two broad categories: Jupiter-family comets (JFCs) and Oort-cloud comets (OCCs). For subgroup definitions and the identification of the individual comets in the plots, we refer to Manfroid et al. (2021) and Hutsemékers et al. (2021). In the plots, we emphasize the measurements of the interstellar comet 2I and of the unusual N2-rich Solar System comet C/2016 R2 (Opitom et al. 2019).
Figure 4 shows Q(NiI)/Q(FeI) as a function of the comet heliocentric distance. No correlation was found between Q(NiI)/Q(FeI) and $r_{h}$ for the Solar System comets (Manfroid et al. 2021). For some comets, Q(NiI)/Q(FeI) was measured at different heliocentric distances, higher than 3 au for C/2000 P1, and it was found to be constant within the error bars. However, comet 3I breaks the rule by showing an extremely high Q(NiI)/Q(FeI) ratio at $r_{h}\simeq$ 2.64 au and a rapid decrease with decreasing $r_{h}$. After one month, at a heliocentric distance of around 2 au, its value became comparable to that of the other Solar System comets. The unexpected behavior of Q(NiI)/Q(FeI) in comet 3I can be clearly seen thanks to the small errors of the measurements. Therefore, we cannot exclude the possibility that other comets exhibited the same behavior, but were simply not regularly monitored with a signal-to-noise ratio good enough to detect it. The two Q(NiI)/Q(FeI) measurements in the interstellar comet 2I seem to show a trend with the heliocentric distance. However, the ratios are equal within the error bars, so the inferred trend is not statistically significant. Therefore, we only considered the average value of Q(NiI)/Q(FeI) for comet 2I in the following.
Figure 5 shows the Q(NiI)/Q(FeI) ratio as a function of the total production rate of the NiI and FeI atoms. On average, OCCs have higher Q(FeI+NiI) rates than JFCs, while the latter show a larger dispersion of the Q(NiI)/Q(FeI) ratio (Hutsemékers et al. 2021). In this plot, comet 3I is clearly exceptional. When the Q(NiI)/Q(FeI) ratio becomes similar to that observed in the other comets, the total Q(FeI+NiI) production rate in comet 3I is one order of magnitude larger than the highest value measured in Solar System comets at heliocentric distance smaller than 1.85 au. Comet 2I, on the other hand, is just ordinary, with a total Q(FeI+NiI) production rate two orders of magnitude smaller than that of comet 3I, at a comparable heliocentric distance. It is also interesting to note that, at $r_{h}\simeq$ 2.75 au, Q(FeI+NiI) in comet C/2016 R2 was a factor of about 2.5 times larger than in comet 3I. However, unlike comet 3I, the metal production in comet C/2016 R2 was dominated by FeI and not NiI. Unfortunately, comet C/2016 R2 was not observed at other heliocentric distances.
Figure 6 shows the correlation between Q(NiI)/Q(FeI) and Q(C2)/Q(CN) previously established for the Solar System comets (Hutsemékers et al. 2021). Given the variation of Q(NiI)/Q(FeI) in comet 3I, one might thus wonder how it fits the relationship. Ignoring the leftmost data point, where Q(CN) temporarily increases faster than Q(C2), $\log$[Q(C2)/Q(CN)] for comet 3I consistently lies between -0.5 and -0.3 at all heliocentric distances (epochs). With $\log$[Q(C2)/Q(CN)] $\lesssim-0.13$ (Bair & Schleicher 2025), comet 3I can thus be confidently classified as a C2-depleted comet, though not as strongly depleted as is reported by Salazar Manzano et al. (2025). Comet 3I perfectly fits the correlation between Q(NiI)/Q(FeI) and Q(C2)/Q(CN) once $\log$[Q(NiI)/Q(FeI)] stabilizes around 0.3 when the comet was at $r_{h}\simeq$ 2 au. While within this correlation, the properties of comet 3I appear very similar to those of comet 2I.
It is also interesting to note that, between $r_{h}$ = 2.64 and 2.19 au, $\log$[Q(CN)/Q(OH)] in comet 3I is in the range of 1.8-1.8 to 2.2-$2.$ These values are consistent with those measured for Solar System comets (A’Hearn et al. 1995), though they are at the higher end. However, our measurements are significantly higher than the ratio of $\log$[Q(CN)/Q(OH)] $\simeq-3.1$ reported by Xing et al. (2025) when comet 3I was at 2.9 au.
In summary, comet 3I exhibits an extremely high Q(NiI)/Q(FeI) ratio at the onset of FeI. Later, this ratio decreases regularly and, at $r_{h}\simeq$ 2 au, becomes indistinguishable from the ratio measured in other Solar System comets and the interstellar comet 2I. On the other hand, the total production rate of NiI and FeI atoms, which was not exceptional at $r_{h}\simeq$ 3 au, strongly increases and, at $r_{h}\simeq$ 2 au, becomes more than one order of magnitude larger than that of other comets observed at comparable or smaller heliocentric distances.
[…]
@@
5 Discussion: The origin of NiI and FeI atoms
The expected blackbody equilibrium temperature, $T$, (in Kelvin) of the cometary surface is approximately given by
$T\simeq 280\;r_{h}^{-1/2}\;,$ (1)